# Dual-Source SPIR over a noiseless MAC without Data Replication or Shared Randomness

### v1

@@ IV-C General case
Suppose that $L_{1}=3$, $L_{2}=4$. $S_{1},S_{2},S_{3}$ are uniformly distributed over $\llbracket 1,2^{nR_{1}/3}\rrbracket$ and $T_{1},T_{2},T_{3},T_{4}$ are uniformly distributed over $\llbracket 1,2^{nR_{2}/2}\rrbracket$, respectively. Server 1 breaks down their files in three parts, i.e., for $i\in\{1,2,3\}$, $F_{1,i}=(F_{1,i,1},F_{1,i,2},F_{1,i,3})$, and Server 2 breaks down their files in two parts, i.e., for $j\in\{1,2,3,4\}$, $F_{2,j}=(F_{2,j,1},F_{2,j,2})$. Servers 1 and 2 engage in a dual-source SPIR protocol from Section IV, interacting with the client $(L_{1}-1)\times(L_{2}-1)=6$ times such that the client obtains one element of each of the six pairs in each of the two following three-by-two matrices,
$M_{1}\triangleq\begin{pmatrix}(F_{1,1,1},S_{1})&(F_{1,2,1}\oplus S_{1},F_{1,3,1}\oplus S_{1})\\ (F_{1,1,2},S_{2})&(F_{1,2,2}\oplus S_{2},F_{1,3,2}\oplus S_{2})\\ (F_{1,1,3},S_{3})&(F_{1,2,3}\oplus S_{3},F_{1,3,3}\oplus S_{3})\end{pmatrix},$
$M_{2}\triangleq\begin{pmatrix}(F_{2,1,1},T_{1})&(F_{2,1,2},T_{3})\\ (F_{2,2,1}\oplus T_{1},T_{1}\oplus T_{2})&(F_{2,2,2}\oplus T_{3},T_{3}\oplus T_{4})\\ (F_{2,3,1}\oplus T_{2},F_{2,4,1}\oplus T_{2})&(F_{2,3,2}\oplus T_{4},F_{2,4,2}\oplus T_{4})\end{pmatrix}.$
Observe that the client can recover part $k\in\{1,2,3\}$ of File $i\in\{1,2,3\}$ from Server 1 with their choices of elements in row $k$ of $M_{1}$. Similarly, the client can recover part $l\in\{1,2\}$ of File $j\in\{1,2,3,4\}$ from Server 2 with their choices of elements in column $l$ of $M_{2}$.
@@
$(L_{1}-2)\times(L_{2}-1)$ sequences $(S_{1,t,i})_{t\in\llbracket 1,L_{1}-2\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}$ uniformly distributed over $\{0,1\}^{nR_{1}/(L_{2}-1)}$, $(L_{2}-2)\times(L_{1}-1)$ sequences $(S_{2,t,j})_{t\in\llbracket 1,L_{2}-2\rrbracket,j\in\llbracket 1,L_{1}-1\rrbracket}$ uniformly distributed over $\{0,1\}^{nR_{2}/(L_{1}-1)}$, the file selection $(Z_{1},Z_{2})\in\mathcal{L}_{1}\times\mathcal{L}_{2}$
2:Server $1$ breaks down File $F_{1,l}$, $l\in\mathcal{L}_{1}$, in $L_{2}-1$ parts denoted by $(F_{1,l,i})_{i\in\llbracket 1,L_{2}-1\rrbracket}$
3:Server $2$ breaks down File $F_{2,l}$, $l\in\mathcal{L}_{2}$, in $L_{1}-1$ parts denoted by $(F_{2,l,j})_{j\in\llbracket 1,L_{1}-1\rrbracket}$
4:Server $1$ forms $(C_{1,t,i})_{t\in\llbracket 1,L_{1}-1\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}$ as follows: for $i\in\llbracket 1,L_{2}-1\rrbracket$,
$(C_{1,1,i}[1],C_{1,1,i}[2])$ $\triangleq(F_{1,1,i},S_{1,1,i})$
$(C_{1,t,i}[1],C_{1,t,i}[2])$ $\triangleq(F_{1,t,i}\oplus S_{1,t-1,i},S_{1,t-1,i}\oplus S_{1,t,i}),\forall t\in\llbracket 2,L_{1}-2\rrbracket$
@@
and for $t\in\llbracket 1,L_{1}-1\rrbracket$, define $C_{1,t,i}\triangleq(C_{1,t,i}[1],C_{1,t,i}[2])$. Consider also the following notation $(\bar{C}_{1,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}\triangleq(C_{1,t,i})_{t\in\llbracket 1,L_{1}-1\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}$
5:Server $2$ forms $(C_{2,t,i})_{t\in\llbracket 1,L_{2}-1\rrbracket,i\in\llbracket 1,L_{1}-1\rrbracket}$ as follows: for $i\in\llbracket 1,L_{1}-1\rrbracket$,
$(C_{2,1,i}[1],C_{2,1,i}[2])$ $\triangleq(F_{2,1,i},S_{2,1,i})$
$(C_{2,t,i}[1],C_{2,t,i}[2])$ $\triangleq(F_{2,t,i}\oplus S_{2,t-1,i},S_{2,t-1,i}\oplus S_{2,t,i}),\forall t\in\llbracket 2,L_{2}-2\rrbracket$
@@
and for $t\in\llbracket 1,L_{2}-1\rrbracket$, define $C_{2,t,i}\triangleq(C_{2,t,i}[1],C_{2,t,i}[2])$. Consider also the following notation $(\bar{C}_{2,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}\triangleq(C_{2,t,i})_{t\in\llbracket 1,L_{2}-1\rrbracket,i\in\llbracket 1,L_{1}-1\rrbracket}$
@@
Consider also the following notation $(Z_{1,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}=(Z_{1,t,i})_{t\in\llbracket 1,L_{1}-1\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}$, $(Z_{2,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}=(Z_{2,t,j})_{t\in\llbracket 1,L_{2}-1\rrbracket,j\in\llbracket 1,L_{1}-1\rrbracket}$
@@
8:     The client and the two servers perform the SPIR protocol in Algorithm 1 with the two sequences $(\bar{C}_{j,k}[1],\bar{C}_{j,k}[2])$ at Server $j\in\llbracket 1,2\rrbracket$ and the selection $(Z_{1,k},Z_{2,k})$ for the client
@@ IV-C1 Server’s privacy with respect to Client
Define $K\triangleq(L_{1}-1)(L_{2}-1)$. Define for $t\in\llbracket 1,K\rrbracket$, $W_{t}\triangleq(\bar{C}_{1,t}[Z_{1,t}],\bar{C}_{2,t}[Z_{2,t}])$, $W_{1:K}\triangleq(\bar{C}_{1,t}[Z_{1,t}],\bar{C}_{2,t}[Z_{2,t}])_{t\in\llbracket 1,K\rrbracket}$, $Y^{n}_{1:K}\triangleq(Y^{n}_{t})_{t\in\llbracket 1,K\rrbracket}$, $\mathbf{A}_{1:K}\triangleq(\mathbf{A}_{t})_{t\in\llbracket 1,K\rrbracket}$, and $(Z_{1,1:K},Z_{2,1:K})\triangleq(Z_{1,t},Z_{2,t})_{t\in\llbracket 1,K\rrbracket})$.
@@
Then, using Lemma 1 and Lemma 2, we prove Server’s privacy with respect to Client as follows. We have
@@
$=I(W_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|F_{1,Z_{1}}F_{2,Z_{2}})+I(Z_{1}Z_{2}Y^{n}_{1:K}\mathbf{A}_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|F_{1,Z_{1}}F_{2,Z_{2}}W_{1:K})$
=$\stackrel{{\scriptstyle(a)}}{{=}}I(Z_{1}Z_{2}Y^{n}_{1:K}\mathbf{A}_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|W_{1:K})$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}I(Z_{1}Z_{2}Y^{n}_{1:K}\mathbf{A}_{1:K};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|W_{1:K})$
$\xrightarrow{n\to\infty}0,$
@@ IV-C2 Client privacy with respect to the Servers
=$\stackrel{{\scriptstyle(a)}}{{=}}I(F^{\prime}_{j,1:K}X_{j,1:K}^{n}\mathbf{A}_{1:K};Z_{1}Z_{2})$
@@
=$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{t=1}^{K}I(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t};Z_{1}Z_{2}Z_{1,t}Z_{2,t}|F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$
=$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{t=1}^{K}I(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t};Z_{1,t}Z_{2,t}|F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$
@@
=$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{t=1}^{K}I(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t};Z_{1,t}Z_{2,t}|S^{\star}F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$
@@
=$\stackrel{{\scriptstyle(g)}}{{=}}\sum_{t=1}^{K}I(S^{\star}\bar{C}_{j,t}[1]\bar{C}_{j,t}[2]X_{j,t}^{n}\mathbf{A}_{t};Z_{1,t}Z_{2,t})$
=$\stackrel{{\scriptstyle(h)}}{{=}}\sum_{t=1}^{K}I(S^{\star};Z_{1,t}Z_{2,t}|\bar{C}_{j,t}[1]\bar{C}_{j,t}[2]X_{j,t}^{n}\mathbf{A}_{t})$
=$\stackrel{{\scriptstyle(i)}}{{=}}0,$
@@
    holds for $k\in\llbracket 1,L_{j}-1\rrbracket$, $l\in\llbracket 1,L_{\bar{j}}-1\rrbracket$, with the definition
    $F^{\prime}_{j,k,l}\triangleq\begin{cases}F_{j,t,l}&\text{ if }k\in\llbracket 1,L_{j}-2\rrbracket\\ (F_{j,L_{j}-1,l},F_{j,L_{j},l})&\text{ if }k=L_{j}-1\end{cases},$
    and $(F^{\prime}_{j,t})_{t\in\llbracket 1,K\rrbracket}\triangleq(F^{\prime}_{j,k,l})_{k\in\llbracket 1,L_{j}-1\rrbracket,l\in\llbracket 1,L_{\bar{j}}-1\rrbracket}$;
@@
    holds with the definition $S^{\star}\triangleq\begin{cases}F^{\prime}_{j,t}\oplus\bar{C}_{j,t}[1]&\text{if }|F^{\prime}_{j,t}|=1\\ F^{\prime}_{j,t}\oplus\bar{C}_{j,t}&\text{if }|F^{\prime}_{j,t}|=2\end{cases}$;
@@
    $=I(S^{\star};\bar{C}_{j,t-1})+I(S^{\star};Z_{1,t}Z_{2,t}F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1}|\bar{C}_{j,t-1})$
@@
    $\leqslant H(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t}|S^{\star})-H(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}S^{\star}F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$

### v2

@@ IV-C General case
Suppose that $L_{1}=3$, $L_{2}=4$. $S_{1},S_{2},S_{3}$ are uniformly distributed over $\llbracket 1,2^{nR_{1}/3}\rrbracket$ and $T_{1},T_{2},T_{3},T_{4}$ are uniformly distributed over $\llbracket 1,2^{nR_{2}/2}\rrbracket$, respectively. Server 1 breaks down their files in three parts, i.e., for $i\in\{1,2,3\}$, $F_{1,i}=(F_{1,i,1},F_{1,i,2},F_{1,i,3})$, and Server 2 breaks down their files in two parts, i.e., for $j\in\{1,2,3,4\}$, $F_{2,j}=(F_{2,j,1},F_{2,j,2})$. Servers 1 and 2 engage in a dual-source SPIR protocol from Section IV-A, interacting with the client $(L_{1}-1)\times(L_{2}-1)=6$ times such that the client obtains one element of each of the six pairs in each of the two three-by-two matrices defined in (11).
$M_{1}\triangleq\begin{pmatrix}(F_{1,1,1},S_{1})&(F_{1,2,1}\oplus S_{1},F_{1,3,1}\oplus S_{1})\\ (F_{1,1,2},S_{2})&(F_{1,2,2}\oplus S_{2},F_{1,3,2}\oplus S_{2})\\ (F_{1,1,3},S_{3})&(F_{1,2,3}\oplus S_{3},F_{1,3,3}\oplus S_{3})\end{pmatrix}\!,\quad M_{2}\triangleq\begin{pmatrix}(F_{2,1,1},T_{1})&(F_{2,1,2},T_{3})\\ (F_{2,2,1}\oplus T_{1},T_{1}\oplus T_{2})&(F_{2,2,2}\oplus T_{3},T_{3}\oplus T_{4})\\ (F_{2,3,1}\oplus T_{2},F_{2,4,1}\oplus T_{2})&(F_{2,3,2}\oplus T_{4},F_{2,4,2}\oplus T_{4})\end{pmatrix}$ (11)
Observe that the client can recover part $k\in\{1,2,3\}$ of File $i\in\{1,2,3\}$ from Server 1 with their choices of elements in row $k$ of $M_{1}$. Similarly, the client can recover part $l\in\{1,2\}$ of File $j\in\{1,2,3,4\}$ from Server 2 with their choices of elements in column $l$ of $M_{2}$. More specifically, the series of requests made by the client to retrieve the desired files is summarized in Table III.
TABLE III: Series of requests for file selection at each server in Example 2.
Server 1 Server 2
Request 1 $F_{1,1,1}$ $S_{1}$ $S_{1}$ $F_{2,1,1}$ $T_{1}$ $T_{1}$ $T_{1}$
Request 2 $F_{1,2,1}\oplus S_{1}$ $F_{1,2,1}\oplus S_{1}$ $F_{1,3,1}\oplus S_{1}$ $F_{2,1,2}$ $T_{3}$ $T_{3}$ $T_{3}$
Request 3 $F_{1,1,2}$ $S_{2}$ $S_{2}$ $F_{2,2,1}\oplus T_{1}$ $F_{2,2,1}\oplus T_{1}$ $T_{1}\oplus T_{2}$ $T_{1}\oplus T_{2}$
Request 4 $F_{1,2,2}\oplus S_{2}$ $F_{1,2,2}\oplus S_{2}$ $F_{1,3,2}\oplus S_{2}$ $F_{2,2,2}\oplus T_{3}$ $F_{2,2,2}\oplus T_{3}$ $T_{3}\oplus T_{4}$ $T_{3}\oplus T_{4}$
Request 5 $F_{1,1,3}$ $S_{3}$ $S_{3}$ $F_{2,3,1}\oplus T_{2}$ $F_{2,3,1}\oplus T_{2}$ $F_{2,3,1}\oplus T_{2}$ $F_{2,4,1}\oplus T_{2}$
Request 6 $F_{1,2,3}\oplus S_{3}$ $F_{1,2,3}\oplus S_{3}$ $F_{1,3,3}\oplus S_{3}$ $F_{2,3,2}\oplus T_{4}$ $F_{2,3,2}\oplus T_{4}$ $F_{2,3,2}\oplus T_{4}$ $F_{2,4,2}\oplus T_{4}$
Recover $F_{1,1}$ $F_{1,2}$ $F_{1,3}$ $F_{2,1}$ $F_{2,2}$ $F_{2,3}$ $F_{2,4}$
Algorithm 3 extends the idea of Example 2 to arbitrary $L_{1}$ and $L_{2}$, using $(L_{1}-1)\times(L_{2}-1)$ sequential interactions between the server and the client. Server $i\in\{1,2\}$ breaks down each file in $L_{\bar{i}}-1$ parts, and each part, except the first, is hidden with random masks. Then, similar to Algorithm 2, in each round $k\in\llbracket 1,(L_{1}-1)\times(L_{2}-1)\rrbracket$, the client selects one of two masked values from Server $i$ via Algorithm 1. Across the $(L_{1}-1)\times(L_{2}-1)$ rounds, the client carefully chooses which values to recover, accumulates them, and combines them with XOR operations to reconstruct the $L_{\bar{i}}-1$ parts of the desired file $F_{i,Z_{i}}$ from Server $i$.
@@
[…]
2:Server 11 breaks down File $F_{1,l}$, $l\in\mathcal{L}_{1}$, in $L_{2}-1$ parts denoted by $(F_{1,l,i})_{i\in\llbracket 1,L_{2}-1\rrbracket}$
3:Server 22 breaks down File $F_{2,l}$, $l\in\mathcal{L}_{2}$, in $L_{1}-1$ parts denoted by $(F_{2,l,j})_{j\in\llbracket 1,L_{1}-1\rrbracket}$
4:Server 11 forms $(C_{1,t,i})_{t\in\llbracket 1,L_{1}-1\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}$ as follows: for $i\in\llbracket 1,L_{2}-1\rrbracket$,
$(C_{1,1,i}[1],C_{1,1,i}[2])$ $\triangleq(F_{1,1,i},S_{1,1,i}),$
$(C_{1,t,i}[1],C_{1,t,i}[2])$ $\triangleq(F_{1,t,i}\oplus S_{1,t-1,i},S_{1,t-1,i}\oplus S_{1,t,i}),\forall t\!\in\!\llbracket 2,L_{1}\!-2\rrbracket,$
@@
and for $t\in\llbracket 1,L_{1}-1\rrbracket$, define $C_{1,t,i}\triangleq(C_{1,t,i}[1],C_{1,t,i}[2]).$ Consider also the following notation
$(\bar{C}_{1,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}\triangleq(C_{1,t,i})_{t\in\llbracket 1,L_{1}-1\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket}.$
5:Server 22 forms $(C_{2,t,i})_{t\in\llbracket 1,L_{2}-1\rrbracket,i\in\llbracket 1,L_{1}-1\rrbracket}$ as follows: for $i\in\llbracket 1,L_{1}-1\rrbracket$,
$(C_{2,1,i}[1],C_{2,1,i}[2])$ $\triangleq(F_{2,1,i},S_{2,1,i}),$
$(C_{2,t,i}[1],C_{2,t,i}[2])$ $\triangleq(F_{2,t,i}\oplus S_{2,t-1,i},S_{2,t-1,i}\oplus S_{2,t,i}),\forall t\!\in\!\llbracket 2,L_{2}-2\rrbracket,$
@@
and for $t\in\llbracket 1,L_{2}-1\rrbracket$, define $C_{2,t,i}\triangleq(C_{2,t,i}[1],C_{2,t,i}[2])$. Consider also the following notation
$(\bar{C}_{2,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}\triangleq(C_{2,t,i})_{t\in\llbracket 1,L_{2}-1\rrbracket,i\in\llbracket 1,L_{1}-1\rrbracket}.$
@@
Consider also the following notation
$(Z_{1,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}\triangleq(Z_{1,t,i})_{t\in\llbracket 1,L_{1}-1\rrbracket,i\in\llbracket 1,L_{2}-1\rrbracket},\quad(Z_{2,k})_{k\in\llbracket 1,(L_{1}-1)(L_{2}-1)\rrbracket}\triangleq(Z_{2,t,j})_{t\in\llbracket 1,L_{2}-1\rrbracket,j\in\llbracket 1,L_{1}-1\rrbracket}$
@@
8:  The client and the two servers perform the SPIR protocol in Algorithm 1 with the two sequences $(\bar{C}_{j,k}[1],\bar{C}_{j,k}[2])$ at Server $j\in\llbracket 1,2\rrbracket$ and the selection $(Z_{1,k},Z_{2,k})$ for the client. The subscript $k$ is used in the notation of the random variables $(X_{1,k}^{n},X_{2,k}^{n},Y_{k}^{n},\mathbf{A}_{k},Z_{1,k},Z_{2,k})$ involved in this SPIR protocol.
@@ IV-C1 Server’s privacy with respect to Client
Define
$K$ $\triangleq(L_{1}-1)(L_{2}-1)$
$W_{t}$ $\triangleq(\bar{C}_{1,t}[Z_{1,t}],\bar{C}_{2,t}[Z_{2,t}]),\forall t\in\llbracket 1,K\rrbracket,$
$W_{1:K}$ $\triangleq(\bar{C}_{1,t}[Z_{1,t}],\bar{C}_{2,t}[Z_{2,t}])_{t\in\llbracket 1,K\rrbracket},$
$Y^{n}_{1:K}$ $\triangleq(Y^{n}_{t})_{t\in\llbracket 1,K\rrbracket},$
$\mathbf{A}_{1:K}$ $\triangleq(\mathbf{A}_{t})_{t\in\llbracket 1,K\rrbracket},$
$(Z_{1,1:K},Z_{2,1:K})$ $\triangleq(Z_{1,t},Z_{2,t})_{t\in\llbracket 1,K\rrbracket}).$
@@
Then, using Lemma 1 and Lemma 2, we prove server’s privacy with respect to the client as follows. We have
@@
$=I(W_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|F_{1,Z_{1}}F_{2,Z_{2}})$
$\phantom{--}+I(Z_{1}Z_{2}Y^{n}_{1:K}\mathbf{A}_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|F_{1,Z_{1}}F_{2,Z_{2}}W_{1:K})$
$\stackrel{{\scriptstyle(a)}}{{=}}I(Z_{1}Z_{2}Y^{n}_{1:K}\mathbf{A}_{1:K};F_{1,\mathcal{L}_{1}}F_{2,\mathcal{L}_{2}}|W_{1:K})$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}I(Z_{1}Z_{2}Y^{n}_{1:K}\mathbf{A}_{1:K};\bar{C}_{1,1:K}\bar{C}_{2,1:K}|W_{1:K})$
$\stackrel{{\scriptstyle(c)}}{{=}}0,$
@@ IV-C2 Client privacy with respect to the servers
$\stackrel{{\scriptstyle(a)}}{{=}}I(F^{\prime}_{j,1:K}X_{j,1:K}^{n}\mathbf{A}_{1:K};Z_{1}Z_{2})$
@@
$\stackrel{{\scriptstyle(b)}}{{=}}\sum_{t=1}^{K}I(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t};Z_{1}Z_{2}Z_{1,t}Z_{2,t}|F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$
$\stackrel{{\scriptstyle(c)}}{{=}}\sum_{t=1}^{K}I(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t};Z_{1,t}Z_{2,t}|F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$
@@
$\stackrel{{\scriptstyle(e)}}{{=}}\sum_{t=1}^{K}I(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t};Z_{1,t}Z_{2,t}|S^{\star}F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$
@@
$\stackrel{{\scriptstyle(g)}}{{=}}\sum_{t=1}^{K}I(S^{\star}\bar{C}_{j,t}[1]\bar{C}_{j,t}[2]X_{j,t}^{n}\mathbf{A}_{t};Z_{1,t}Z_{2,t})$
$\stackrel{{\scriptstyle(h)}}{{=}}\sum_{t=1}^{K}I(S^{\star};Z_{1,t}Z_{2,t}|\bar{C}_{j,t}[1]\bar{C}_{j,t}[2]X_{j,t}^{n}\mathbf{A}_{t})$
$\stackrel{{\scriptstyle(i)}}{{=}}0,$
@@
    holds for $k\in\llbracket 1,L_{j}-1\rrbracket$, $l\in\llbracket 1,L_{\bar{j}}-1\rrbracket$, with the definition F’_j, k,l ≜{Fj, t,l if k ∈⟦1, Lj-2 ⟧(Fj,Lj-1,l,Fj,Lj,l) if k = Lj-1, and $(F^{\prime}_{j,t})_{t\in\llbracket 1,K\rrbracket}\triangleq(F^{\prime}_{j,k,l})_{k\in\llbracket 1,L_{j}-1\rrbracket,l\in\llbracket 1,L_{\bar{j}}-1\rrbracket}$;
@@
    holds with the definition S^⋆≜{F’j,t⊕¯Cj,t[1] if —F’j,t—=1F’j,t⊕¯Cj,tif —F’j,t—=2;
@@
    $=I(S^{\star};\bar{C}_{j,t-1})$
    $\phantom{-}+I(S^{\star};Z_{1,t}Z_{2,t}F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1}|\bar{C}_{j,t-1})$
@@
    $\leqslant H(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t}|S^{\star})$
    $\phantom{-}-H(F^{\prime}_{j,t}X_{j,t}^{n}\mathbf{A}_{t}|Z_{1,t}Z_{2,t}S^{\star}F^{\prime}_{j,1:t-1}X_{j,1:t-1}^{n}\mathbf{A}_{1:t-1})$